OpenStudy (anonymous):

How would I find out which the equation of the line passing through the point (6, -2) and parallel to y = x + 7 is? I find this stuff hard to understand:/

OpenStudy (anonymous):

What is parallel to 1x?

OpenStudy (aroub):

Read this: http://openstudy.com/#/updates/4efa3e57e4b01ad20b51b35d it might help ( some how ) :D

OpenStudy (aroub):

Burn?

OpenStudy (anonymous):

I'm still confused..

OpenStudy (aroub):

What's confusing you?

OpenStudy (anonymous):

How you got the other problem

OpenStudy (aroub):

lets try to do this problem ok?

OpenStudy (anonymous):

Okay

OpenStudy (aroub):

you know that y=mx+b where m is the slope and b is the y-intercept right?

OpenStudy (anonymous):

Right.

OpenStudy (aroub):

Parallel= same slope -ALWAYS!

OpenStudy (aroub):

y = x + 7 what's the slope here?

OpenStudy (anonymous):

x?

OpenStudy (aroub):

No, well right this is a bit tricky... it's actually 1 because 1x is the same as x so they don't bother writing 1

OpenStudy (anonymous):

Oh, it is?

OpenStudy (aroub):

Yup :)

OpenStudy (anonymous):

oh..

OpenStudy (aroub):

btw, which part you didn't understand in the previous question? because it's exactly like this question...

OpenStudy (anonymous):

I think it's after you started to substitute everything

OpenStudy (aroub):

oh I substituted because you have to find the y-intercept

OpenStudy (anonymous):

so one sec. Right now would it be y= 1+ 7?

OpenStudy (aroub):

Meverett?

OpenStudy (aroub):

wait, what did you do??

OpenStudy (anonymous):

I just confused myself even more://

OpenStudy (aroub):

I would love to help more! but I sooo have to go now! I'm so sorry :(

OpenStudy (anonymous):

It's okay, thank you though!

OpenStudy (karatechopper):

hmm mind if i help?

OpenStudy (aroub):

thank you karate :D

OpenStudy (anonymous):

I'd love it if you could help!

OpenStudy (karatechopper):

alright

OpenStudy (karatechopper):

ok so lets start with this. do you know how to write and equation of a line if i gave u a point and slope?

OpenStudy (anonymous):

Possibly..

OpenStudy (karatechopper):

want to try one? tht will warm u up for it

OpenStudy (anonymous):

Yes please

OpenStudy (karatechopper):

if i give u the point (3,1) with a slope of 1/3 write this as an equation in the form of slope intercept form

OpenStudy (anonymous):

1=1/3 + 3? I think that's completely wrong

OpenStudy (karatechopper):

umm it is.. sorry:)

OpenStudy (karatechopper):

lets learn to solve tht one first

OpenStudy (karatechopper):

k?

OpenStudy (anonymous):

okay

OpenStudy (karatechopper):

so lets look back at my question to u: if i give u the point (3,1) with a slope of 1/3 write this as an equation in the form of slope intercept form

OpenStudy (karatechopper):

do you know the equation of slope intercept form?

OpenStudy (anonymous):

isn't it y=mx+b

OpenStudy (karatechopper):

correct!

OpenStudy (karatechopper):

ok now can u tell me what the variable is for slope in that equation?

OpenStudy (anonymous):

So you said 1/3 is the slope, so y would be 1 and b would be 3? :// i'm not sure

OpenStudy (karatechopper):

no.

OpenStudy (karatechopper):

u got the slope part right. of we were to write the equation containing slope only. then it would look like this: y=1/3x+b <---- see where i am going?

OpenStudy (anonymous):

yeah

OpenStudy (karatechopper):

so if we keep this equation.. we have to have the y intercept and slope. so in order for us to get the y intercept, we must solve for it. do u know what we do to solve for y int?

OpenStudy (anonymous):

no

OpenStudy (karatechopper):

we plug in the x and y values like this. we use the point we are given. 1=1/3(3)+b

OpenStudy (karatechopper):

now can u solve for b?

OpenStudy (anonymous):

well 1/3(3) is 1, right?

OpenStudy (karatechopper):

yes

OpenStudy (anonymous):

actually, I'm lost lol

OpenStudy (karatechopper):

haha ok

OpenStudy (karatechopper):

i just asked u to solve for b 1=1/3(3)+b

OpenStudy (anonymous):

how would I do that?

OpenStudy (karatechopper):

have u solved an equation before?

OpenStudy (anonymous):

Yes, but I am terrible in math.

OpenStudy (karatechopper):

its fine, i should have said solve the equation

OpenStudy (anonymous):

I am sorry, but I have to go eat, so i'll be back in a bit. thank you though

OpenStudy (karatechopper):

yes

OpenStudy (karatechopper):

your welcome